Junior's boat will go 15 miles per hour in still water. If he can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current?
Found 2 solutions by Nate, mathdude2:
Answer by Nate(3500)

9/15 = time = 0.6 hours
12/x = 0.6 rate in current
12 = 0.6x
20 = x
Rate with Current: 20mph
Rate without Current: 15mph
20 - 15 = 5
Current's Speed: 5mph

Answer by mathdude2(1)


The above solution is not correct.
Here is the correct solution:

Using the formula t=s/v
where t=time, s=distance, v=speed
12/(15+x)=9/(15-x)

Solving for X, multiply both sides by (15-x)/(15+x)

12(15-x)=9(15+x)

180-12x=135+9x

180-135=12x+9x

45=21x

x=45/21=15/7 = 2 1/7

Then the speed of the current x=15/7 or 2 1/7 miles per hour.